a ball is thrown at an angle theta to the ground,when the wind exerts a force mg on it,state with reasons whether it will be a parabola or not???if so state how will be its axis of symmetry???!

Grade:

1 Answers

vikas askiitian expert
13 years ago

motion will be straight line y=x is the equation of its motion

prof

      initial velocity of particle is ucos@i +usin@j

      total accleration is a=-gi-gj

      applying s(x,y)=ut+(at^2)/2

      xi+yj= (ucos@i+usin@j)t +-(gi+gj)t^2/2

       x=ucos@t-gt^2/2.....1   ,   y=usin@t-gt^2/2.......2

       x-y=(ucos@-usin@)t........3

 putting 3 in 1 or 2 we get (x-y)=0 and tan@=y/x

        we get the eq of path as straight line